# MVGS: Multi-view-regulated Gaussian Splatting for Novel View Synthesis

### v2

@@ 1 Introduction
Photorealistic rendering towards unbounded scenes or single object is proven to hold significant values in both the industrial and academic areas, e.g. multi-media generation, virtual reality, and autonomous driving. Conventional primitive-based representations such as mesh and point cloud (Botsch et al., 2005; Lassner & Zollhofer, 2021; Yifan et al., 2019; Munkberg et al., 2022) make real-time rendering possible with the help of efficient rasterization. Although such a rendering mechanism delivers high efficiency, revealing a fine-grained and sufficiently precise appearance is still struggling, where blurry artifacts and discontinuity happen. On the contrary, implicit representation (Erler et al., 2020) and neural radiance field (Mildenhall et al., 2021; Barron et al., 2022; Müller et al., 2022)(NeRF), employ the multi-layer perceptron (MLP) to improve the potentiality of rendering high-fidelity geometries where more detailed structures are preserved. However, their inference efficiency is still limited even with accelerating operators such as Instant-NGP (Müller et al., 2022).
@@ 2 Related Work
Significant advancements have been achieved in novel-view synthesis, particularly since the introduction of NeRF (Barron et al., 2021; Mildenhall et al., 2021), which employs MLP to parameterize the geometry and view-dependent appearance with the help of implicitly defined radiance field. Moreover, the training and inference efficiency of NeRF has been enhanced using hash-grid (Müller et al., 2022) and explicitly defined samplers (Li et al., 2023). Built on top of the radiance field, NeuS (Wang et al., 2021), NeuS2 (Wang et al., 2023), and HF-NeuS (Wang et al., 2022) also perform more precise surface reconstruction against traditional MVS fusion such as MeshMVS (Shrestha et al., 2021). Given all the advantages of neural rendering, its efficiency is still not implausible. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged, demonstrating impressive real-time NVS performance. Gaussian-based methods, typically represented by 3D Gaussian Splating (Kerbl et al., 2023), are recent advancements in 3D reconstruction, enabling high-quality and real-time rendering.
@@ 3 Methodology
Gaussian Splatting (Kerbl et al., 2023) is recently proposed for real-time novel-view synthesis and high-fidelity 3D geometric reconstruction. Instead of employing implicit representations such as density field in NeRF(Mildenhall et al., 2021) and SDF in NeuS (Wang et al., 2021), Gaussian Splatting leverages a set of anisotropic 3D Gaussians comprising their locations, colors, covariances, and opacities to parameterize a scene. Such an explicit representation dramatically improves the training and inference efficiency compared to previous methods like NeRF and NeuS. In the rendering process, Gaussian Splatting also adopts the point-based volume rendering technique (Kopanas et al., 2021; 2022a) following NeRF. As denoted in Fig. 2 (a), we diagram that NeRF cannot receive multi-view supervision in a training iteration due to its point-sampling strategy and implicit representation. The view-dependent radiance $\mathbf{C}(.)$ of each pixel $p$ in the image with the camera extrinsics $E$ and intrinsics $K$ is calculated by blending a set of 3D Gaussians along ray $r(p,E,K)$. While NeRF (Mildenhall et al., 2021) approximately blends with points assigned by a sampler (Li et al., 2023) in the radiance field, 3DGS precisely blends by rasterizing with NN parameterized kernels $\mathbf{G}(r)=\{g_{i}~|~i=1,\dots,N\}$ along ray $r(p,E,K)$. Assuming that the color $c_{i}\in\mathbb{R}^{3}$, the opacity $o_{i}\in\mathbb{R}$, and the covariance $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ describe the attributes of the $i$-th Gaussian $g_{i}$ respectively, the rendered pixel radiance $\mathbf{C}(r)$ is represented as
@@
where the color $c_{i}$ is weighted by the transmittance $\alpha_{i}=o_{i}\exp{({-\frac{1}{2}(x_{i})^{T}\Sigma_{i}^{-1}(x_{i})})}$. Here $x_{i}$ denotes the distance between the position $\mu_{i}\in\mathbb{R}^{3}$ of Gaussian kernel and the query pixel pp. NN represents the number of 3D Gaussians. Considering that 3DGS directly blends with the individually parameterized kernels $g_{i}=\{o_{i},c_{i},\Sigma_{i}\}$, which makes it possible to investigate optimization towards specific kernels across different views, we propose to improve the performance of 3DGS by a novel multi-view constraint in Sec. 3.1, stabilize rays by enriching different intrinsic setups in Sec. 3.2, and adapt the densification strategy in Sec. 3.3 and Sec. 3.4.
@@ 3.4 Multi-view Augmented Densification
where $H(\cdot)$ is Heaviside function, returning 11 if the input is larger or equal 0. $\tau$ is a predefined hyperparameter. In this way, when the discrepancies between each view become large, the extent of 3D Gaussian densification is also enhanced. Consequently, our proposed multi-view augmented densification strategy allows 3D Gaussians to fit better for each view and capture finer details.
@@ 4 Experiments
HexPlane-Slim (Cao & Johnson, 2023) 31.04 0.973 0.044 2.5 38

### v3

@@ 1 Introduction
Photorealistic rendering towards unbounded scenes or single object is proven to hold significant values in both the industrial and academic areas, e.g. multi-media generation, virtual reality, and autonomous driving. Conventional primitive-based representations such as mesh and point cloud (Botsch et al., 2005; Lassner and Zollhofer, 2021; Yifan et al., 2019; Munkberg et al., 2022) make real-time rendering possible with the help of efficient rasterization. Although such a rendering mechanism delivers high efficiency, revealing a fine-grained and sufficiently precise appearance is still struggling, where blurry artifacts and discontinuity happen. On the contrary, implicit representation (Erler et al., 2020) and neural radiance field (Mildenhall et al., 2021; Barron et al., 2022; Müller et al., 2022a)(NeRF), employ the multi-layer perceptron (MLP) to improve the potentiality of rendering high-fidelity geometries where more detailed structures are preserved. However, their inference efficiency is still limited even with accelerating operators such as Instant-NGP (Müller et al., 2022b).
@@ 2 Related Work
Significant advancements have been achieved in novel-view synthesis, particularly since the introduction of NeRF (Barron et al., 2021; Mildenhall et al., 2021), which employs MLP to parameterize the geometry and view-dependent appearance with the help of implicitly defined radiance field. Moreover, the training and inference efficiency of NeRF has been enhanced using hash-grid (Müller et al., 2022a) and explicitly defined samplers (Li et al., 2023). Built on top of the radiance field, NeuS (Wang et al., 2021), NeuS2 (Wang et al., 2023), and HF-NeuS (Wang et al., 2022) also perform more precise surface reconstruction against traditional MVS fusion such as MeshMVS (Shrestha et al., 2021). Given all the advantages of neural rendering, its efficiency is still not implausible. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged, demonstrating impressive real-time NVS performance. Gaussian-based methods, typically represented by 3D Gaussian Splating (Kerbl et al., 2023), are recent advancements in 3D reconstruction, enabling high-quality and real-time rendering.
@@ 3 Methodology
Gaussian Splatting (Kerbl et al., 2023) is recently proposed for real-time novel-view synthesis and high-fidelity 3D geometric reconstruction. Instead of employing implicit representations such as density field in NeRF(Mildenhall et al., 2021) and SDF in NeuS (Wang et al., 2021), Gaussian Splatting leverages a set of anisotropic 3D Gaussians comprising their locations, colors, covariances, and opacities to parameterize a scene. Such an explicit representation dramatically improves the training and inference efficiency compared to previous methods like NeRF and NeuS. In the rendering process, Gaussian Splatting also adopts the point-based volume rendering technique (Kopanas et al., 2021; 2022a) following NeRF. As denoted in Fig. 2 (a), we diagram that NeRF cannot receive multi-view supervision in a training iteration due to its point-sampling strategy and implicit representation. The view-dependent radiance $\mathbf{C}(.)$ of each pixel $p$ in the image with the camera extrinsics $E$ and intrinsics $K$ is calculated by blending a set of 3D Gaussians along ray $r(p,E,K)$. While NeRF (Mildenhall et al., 2021) approximately blends with points assigned by a sampler (Li et al., 2023) in the radiance field, 3DGS precisely blends by rasterizing with $N$ parameterized kernels $\mathbf{G}(r)=\{g_{i}~|~i=1,\dots,N\}$ along ray $r(p,E,K)$. Assuming that the color $c_{i}\in\mathbb{R}^{3}$, the opacity $o_{i}\in\mathbb{R}$, and the covariance $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ describe the attributes of the $i$-th Gaussian $g_{i}$ respectively, the rendered pixel radiance $\mathbf{C}(r)$ is represented as
@@
where the color $c_{i}$ is weighted by the transmittance $\alpha_{i}=o_{i}\exp{({-\frac{1}{2}(x_{i})^{T}\Sigma_{i}^{-1}(x_{i})})}$. Here $x_{i}$ denotes the distance between the position $\mu_{i}\in\mathbb{R}^{3}$ of Gaussian kernel and the query pixel $p$. NN represents the number of 3D Gaussians. Considering that 3DGS directly blends with the individually parameterized kernels $g_{i}=\{o_{i},c_{i},\Sigma_{i}\}$, which makes it possible to investigate optimization towards specific kernels across different views, we propose to improve the performance of 3DGS by a novel multi-view constraint in Sec. 3.1, stabilize rays by enriching different intrinsic setups in Sec. 3.2, and adapt the densification strategy in Sec. 3.3 and Sec. 3.4.
@@ 3.4 Multi-view Augmented Densification
where $H(\cdot)$ is Heaviside function, returning $1$ if the input is larger or equal 0. $\tau$ is a predefined hyperparameter. In this way, when the discrepancies between each view become large, the extent of 3D Gaussian densification is also enhanced. Consequently, our proposed multi-view augmented densification strategy allows 3D Gaussians to fit better for each view and capture finer details.
@@ 4 Experiments
HexPlane-Slim (Cao and Johnson, 2023) 31.04 0.973 0.044 2.5 38